OpenStudy (anonymous):

(x-1)*+(2x+3)*=-10x-10

hartnn (hartnn):

@sapo Hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) can you post the question more clearly ? (x-1)* doesn't make much sense, whats the exponent ?

OpenStudy (anonymous):

dis the question but these one r square 2

OpenStudy (anonymous):

i mean *these one

OpenStudy (anonymous):

(x-1)*+(2x+3)*=-10x-10

hartnn (hartnn):

is the question, \((x-1)^2+(2x+3)^2=-10x-10\) ??

OpenStudy (anonymous):

yes

hartnn (hartnn):

use \((a+b)^2 = a^2+2ab+b^2\) and \((a-b)^2 = a^2-2ab+b^2\) so what about (x-1)^2 ?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

in this question i want to from where+9 and12x and -2x came

hartnn (hartnn):